OpenStudy (stormoshawty):

Solve this problem using elimination. 1/3x + 1/2y = 0 1/2x + 1/5y = 11/5

OpenStudy (owlcoffee):

Fractions are a little tricky, but once you get used to them, they are like any number, so let's see: \[\frac{ 1 }{ 3 }x+\frac{ 1 }{ 2 }y=0\] \[\frac{ 1 }{ 2 }x+\frac{ 1 }{ 5 }y=\frac{ 11 }{ 5 }\] Let's make the variables opposite in this problem, and use elimination, let's begin by multiplying the first equation by -6 and the second by 4, operate and order up: \[-2x-3y=0\] \[2x+\frac{ 4 }{ 5 }y=\frac{ 44 }{ 5 }\] So, sum up both equations, to get rid of the x term and associating the y terms we get: \[(\frac{ 4 }{ 5 }-3)y=\frac{ 44 }{ 5 }\] Solving the operation on the inside of the parenthessis: \[(\frac{ 4-15 }{ 5 })y=\frac{ 44 }{ 5 }\] operating a little and reordering: \[-11y=44\] And we have found that y=(-4) Let's take the initial equation (any of them) and replace and find the x value: \[\frac{ 1 }{ 3 }x+\frac{ 1 }{ 2 }(-4)=0\] multiplying and ordering up, also adding 2 on both sides: \[\frac{ 1 }{ 3 }x=2\] solving for x we get that x=6, so we conclude that the solution is: \[x=6\] \[y=-4\]